# Heterophily-Aware Fair Recommendation using Graph Convolutional Networks

### v1

@@ 5.4 Hyper-parameters study
To study the impact of the number of GNN layers, we select two datasets Beauty and Epinions, and investigate the potential effect of over-smoothing over them. We evaluate the impact of the number of layers by ranging them from 1 to 6, and use NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 4 and 5. For the Epinions dataset, our model outperforms all the baselines across all the metrics. In all the methods, NDCG increases until reaching the optimal point and then starts to decline due to the over-smoothing phenomenon. As the number of layers increases, PRU and PRI metrics decrease, indicating that increasing the number of layers up to a certain point can alleviate the unfairness caused by popularity bias. The effect of over-smoothing becomes less noticeable due to the use of the weighted sum approach for the final node representation. On the other hand in the Beauty dataset, although we use the same weighted sum approach, the effect of the over-smoothing problem is evident from the initial stages of the models. The performance of the LightGCN model drastically declines after 4 GNN layers. The performance of HetroFair decreases too, but with a smaller slope and it still outperforms all the baselines over the Beauty dataset. This shows its robust performance across different settings.
@@
The embeddings dimension is an important factor in graph-based recommendation systems, representing the latent factors size of users and items. We consider two datasets CDs and Epinions, and compare the performance of the models across three embeddings sizes: $64$ (similar to LightGCN and r-AdjNorm), $128$ (the default value for HetroFair) and $256$ (the used value by APDA). We utilize NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 2 and 3. As can be seen in the figures, over both datasets and by increasing the embeddings dimension, NDCG of HetroFair increases, too. In other words, in our model NDCG has positive correlation with the embeddings dimension. For all values of the embeddings dimension, HetroFair outperforms the other methods, in terms of both accuracy and fairness measures. It can also be seen in the figures that the dependence of PRI on the dimension size is negligible. In Figures 1(a) and 2(a), the plots exhibit steeper slopes when the dimensions change from $64$ to $128$, compared to the cases where the dimensions change from $128$ to $264$. This means that the rates of accuracy changes decrease from 128 dimensions and beyond. According to our time complexity analysis in Section 4, time complexity of HetroFair increases linearly with the dimension of the embeddings. Therefore, considering $128$ as the default dimension for embeddings seems to be a good trade-off between accuracy and complexity. In the investigation of the baseline models, it can be pointed out that no model consistently outperforms the other baselines and by revising the parameters, the relative performances of the models change with respect to each other. For example in Figure 3, we see that in most cases, LightGCN achieves better fairness scores than APDA, suggesting that the simple LightGCN method effectively captures data characteristics, particularly when a larger embeddings dimension is used. On the other hand, r-AdjNorm has the worst performance in fairness metrics, indicating that it fails to make a trade-off between accuracy and fairness.
@@
5.5 Short-head and long-tail items performance
To evaluate the effectiveness of HetroFair in recommending long-tail items, we modify the prior test set and construct two following test sets, one for long-tail items and the other for short-head items. For the long-tail test set, we exclude all short-head items from the $\tilde{O}_{u}^{+}$ sets of all users $u$. For the short-head test set, we remove all long-tail items existing in the $\tilde{O}_{u}^{+}$ sets. We sort the items based on their degree in descending order and consider the top $20\%$ of them as short-head items, treating the rest as long-tail items. Figure 8 depicts the results of this investigation, over all the six datasets. These figures illustrate that for the category of long-tail items, in most cases HetroFair achieves better accuracies. This is compatible with the results presented in Table 3, which indicate a reduction in the fairness metrics. Moreover, this improvement coincides with an increase in accuracy in the category of short-head items. Combining the results of Table 3 with the results presented in Figure 8 depicts that our performance improvement is not solely tied to long-tail items. The simultaneous improvement of HetroFair in recommending short-head items also plays a substantial role in reinforcing its overall performance.

### v2

@@ 5.4 Hyper-parameters study
To study the impact of the number of GNN layers, we select two datasets Beauty and Epinions, and investigate the potential effect of over-smoothing over them. We evaluate the impact of the number of layers by ranging them from 1 to 6, and use NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 2 and 3. For the Epinions dataset, our model outperforms all the baselines across all the metrics. In all the methods, NDCG increases until reaching the optimal point and then starts to decline due to the over-smoothing phenomenon. As the number of layers increases, PRU and PRI metrics decrease, indicating that increasing the number of layers up to a certain point can alleviate the unfairness caused by popularity bias. The effect of over-smoothing becomes less noticeable due to the use of the weighted sum approach for the final node representation. On the other hand in the Beauty dataset, although we use the same weighted sum approach, the effect of the over-smoothing problem is evident from the initial stages of the models. The performance of the LightGCN and PC models drastically declines after 5 GNN layers. The performance of HetroFair decreases too, but with a smaller slope and it still outperforms all the baselines over the Beauty dataset. This shows its robust performance across different settings.
@@
The embeddings dimension is an important factor in graph-based recommendation systems, representing the latent factors size of users and items. We consider two datasets CDs and Epinions, and compare the performance of the models across three embeddings sizes: $64$ (similar to LightGCN and r-AdjNorm), $128$ (the default value for HetroFair) and $256$ (the used value by APDA). We utilize NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 4 and 5. As can be seen in the figures, over both datasets and by increasing the embeddings dimension, NDCG of HetroFair increases, too. In other words, in our model NDCG has positive correlation with the embeddings dimension. For all values of the embeddings dimension, HetroFair outperforms the other methods, in terms of both accuracy and fairness measures. It can also be seen in the figures that the dependence of PRI on the dimension size is negligible. Considering the NDCG metric in Figures 4 and 5, the plots exhibit steeper slopes when the dimensions change from $64$ to $128$, compared to the cases where the dimensions change from $128$ to $264$. This means that the rates of accuracy changes decrease from 128 dimensions and beyond. According to our time complexity analysis in Section 4, time complexity of HetroFair increases linearly with the dimension of the embeddings. Therefore, considering $128$ as the default dimension for embeddings seems to be a good trade-off between accuracy and complexity.
In the investigation of the baseline models, it can be pointed out that no model consistently outperforms the other baselines and by revising the parameters, the relative performances of the models change with respect to each other. For example in Figure 4, we see that in most cases, LightGCN achieves better fairness scores than APDA, suggesting that the simple LightGCN method effectively captures data characteristics, particularly when a larger embeddings dimension is used. On the other hand, r-AdjNorm has the worst performance in fairness metrics, indicating that it fails to make a trade-off between accuracy and fairness.
@@
5.5 Short-head and long-tail items performance
To evaluate the effectiveness of HetroFair in recommending long-tail items, we modify the prior test set and construct two following test sets, one for long-tail items and the other for short-head items. For the long-tail test set, we exclude all short-head items from the $\tilde{O}_{u}^{+}$ sets of all users $u$. For the short-head test set, we remove all long-tail items existing in the $\tilde{O}_{u}^{+}$ sets. We sort the items based on their degree in descending order and consider the top $20\%$ of them as short-head items, treating the rest as long-tail items.
Figure 8 depicts the results of this investigation, over all the six datasets. These figures illustrate that for the category of long-tail items, in most cases HetroFair achieves better accuracies. This is compatible with the results presented in Table 3, which indicate a reduction in the fairness metrics. Moreover, this improvement coincides with an increase in accuracy in the category of short-head items. Combining the results of Table 3 with the results presented in Figure 8 depicts that our performance improvement is not solely tied to long-tail items. The simultaneous improvement of HetroFair in recommending short-head items also plays a substantial role in reinforcing its overall performance.